A ball of clay with a mass of 0.250 kg is travelin
Answer #1

Given: mass of ball m =0.250 kg, speed of ball v= 12 m/s , Irod= 0.800 kg.m2, L=1.20m

Irod = 1/12 ML2

Mass of rod M= (12×0.800)/1.202 = 6.67 kg

Angular momentum conservation equation about rod's pivot point is:

Li = Lf =mvr= Iball+rod\omega

Note, r = L/2 = 0.6 m

Iball+rod = mr2 +Irod = .89 kg.m2

If vf is the final velocity of clay ball, \omega = vf/r .

Angular momentum L = mvr = Iball+rod\omega

\omega = mvr /Iball+rod = (0.250×12×0.6)/0.89

  \omega = 2.02 rad/s and vf = 2.02×0.6 = 1.21 m/s